At a local grocery store, three bags of peanuts and four bags of almonds cost $27.45. At the same store, five bags of peanuts and two bags of almonds cost $24.05. If Nyesha wants to buy one bag of peanuts and one bag of almonds, how much money does she need?

Answer by checkley77(12844)

3p+4a=27.45
5p+2a=24.05 mutiply by -2 & add
-10p-4a=-48.10
3p+4a=27.45
----------------------
-7p=-20.65
p=-20.65/-7
p=2.95 cost of the peanuts.
3*2.95+4a=27.45
8.85+4a=27.45
4a=27.45-8.85
4p=18.60p=18.60/4
p=4.65 for the almonds.
2.95+4.65=7.60 cost of one bag of peanuts + one bag of almonds.